### v1

@@ Discussion
CIDEr
Consensus-based Image Description Evaluation (CIDEr) [57] is a metric designed to assess the similarity of a generated text to multiple reference texts, emphasizing consensus. It is used for medical report generation to evaluate the relevance and quality of the generated content.
Clinical efficacy
Clinical Efficacy measures the practical effectiveness of a generated medical report in a clinical setting, focusing on its impact on clinical outcomes and decision-making. It includes precision, recall, and F1, by annotating the generated report and comparing it with the ground truth report via 14 disease classification labels in CheXbert [58].
Micro precision calculates the precision of the model by considering all true positives and false positives across all classes, providing a detailed view of the model’s performance:
Micro $\text{Micro precision}=\frac{\sum\text{TP}}{\sum\text{TP}+\sum\text{FP}}$
Micro recall aggregates the true positives and false negatives across all classes to compute recall, offering a comprehensive assessment of the model’s ability to identify relevant instances:
Micro $\text{Micro recall}=\frac{\sum\text{TP}}{\sum\text{TP}+\sum\text{FN}}$
Micro F1 is the harmonic mean of Micro Precision and Micro Recall, providing a balanced measure of the model’s accuracy and completeness across all classes:
@@
The dataset used for pre-training is described in Section Method. The datasets used for evaluation are listed in Extended Data Table LABEL:tab:datasets.
@@
The implementation of ConceptCLIP will be available at https://github.com/JerrryNie/ConceptCLIP. The weights of ConceptCLIP will be released at https://huggingface.co/JerrryNie/ConceptCLIP.
@@
Y.N., S.H., Y.B., and H.C. conceived and designed the work. Y.N., S.H., Y.B., Y.W., and Z.C. collected the data for pre-training and downstream task evaluation. Y.N. and S.H. contributed to the pre-training technical implementation. Y.B. contributed to the technical implementation of explainable AI. Y.N. and S.H. evaluated medical image diagnosis tasks. Y.N. evaluated cross-modal retrieval tasks. Y.B. evaluated explainable AI tasks. Y.W. evaluated whole-slide image analysis tasks. Z.C. evaluated medical report generation tasks. S.Y. evaluated medical visual question answering tasks. Y.N., Y.B., and S.H. wrote the manuscript with inputs from all authors. All authors reviewed and approved the final paper. H.C. supervised the research.
@@ Appendix A Extended Data
Table 1: Medical image analysis datasets.
Dataset Website Description
@@
SIIM-ACR [59] https://www.kaggle.com/c/siim-acr-pneumothorax-segmentation A binary classification dataset containing chest radiographic images.
Covid-CXR2 [60] https://www.kaggle.com/datasets/andyczhao/covidx-cxr2 A dataset of over 16,000 chest X-ray images from more than 15,100 patients across 51 countries, including 2,300 positive COVID-19 images, designed to differentiate between no pneumonia, non-COVID-19 pneumonia, and COVID-19 pneumonia in the COVIDx V8A dataset, while the COVIDx V8B dataset focuses on detecting COVID-19 positive and negative cases.
NLM-TB [25] https://www.ncbi.nlm.nih.gov/pmc/articles/PMC4256233/ A radiology image dataset with 800 total images with the label of normal or tuberculosis.
LC25000 (Colon) [61] https://github.com/tampapath/lung_colon_image_set A dataset comprises 10,000 pathology images from colon tissue, featuring benign tissues and adenocarcinomas.
UBIBC [62] https://www.kaggle.com/datasets/vuppalaadithyasairam/ultrasound-breast-images-for-breast-cancer This dataset consists of ultrasound images related to breast cancer, both benign and malignant.
PCam200 [63] https://github.com/enigmanx20/patchtcga A dataset of pathological H&E images created from the Camelyon2016 challenge dataset [64].
RSNA [48] https://www.rsna.org/rsnai/ai-image-challenge/rsna-pneumonia-detection-challenge-2018 A dataset of about 30,000 frontal view chest X-ray images, each labeled for binary classification to indicate the presence of pneumonia.
Brain Tumor CT [65] https://www.kaggle.com/datasets/murtozalikhon/brain-tumor-multimodal-image-ct-and-mri A dataset of CT scans for brain tumor detection and analysis. It features high-resolution images from multiple patients, labeled with tumor types (e.g., glioma, meningioma) and their locations in the brain, designed to aid in developing AI models for automatic detection, classification, and segmentation of brain tumors.
Brain Tumor MRI [65] https://www.kaggle.com/datasets/murtozalikhon/brain-tumor-multimodal-image-ct-and-mri A dataset of MRI scans, for brain tumor detection and analysis. It features high-resolution images from multiple patients, labeled with tumor types (e.g., glioma, meningioma) and their locations in the brain, designed to aid in developing AI models for automatic detection, classification, and segmentation of brain tumors.
DDSM [66] https://www.kaggle.com/datasets/skooch/ddsm-mammography A dataset containing 55,890 pre-processed images from the DDSM [67] and CBIS-DDSM [68] datasets, formatted as 299x299 pixel tfrecords for TensorFlow, with an imbalance of 14% positive and 86% negative examples, incorrectly split into test and validation sets.
Breast Cancer [69] https://www.kaggle.com/datasets/hayder17/breast-cancer-detection/ A dataset of 3,383 annotated mammogram images focused on breast tumors, exported from Roboflow, designed for building and testing deep learning models for tumor detection.
@@
RFMiD2 [70] https://www.mdpi.com/2306-5729/8/2/29 A multi-label dataset of fundus images annotated by three eye specialists.
MedFMC (Colon) [71] https://github.com/openmedlab/MedFM A dataset focused on facilitating the detection of early-stage cancer cells in tissue slides, enabling pathologists to classify and quantify cancerous regions for lesion and non-lesion classes.
VinDr-Mammo [72] https://vindr.ai/datasets/mammo A large-scale benchmark dataset for computer-aided detection and diagnosis in full-field digital mammography.
VinDr-PCXR [50] https://vindr.ai/datasets/pediatric-chest-x-ray A dataset focuses on pediatric chest X-rays for the interpretation of common thoracic diseases.
VinDr-CXR [49] https://vindr.ai/datasets/cxr A dataset of chest x-ray images containing 18,000 high-quality postero-anterior scans with detailed localization of 22 critical findings and classification of 6 thoracic diseases, annotated by experienced radiologists from major hospitals in Vietnam. To make the official training and testing label consistent, we filter out part of inconsistent training data, thus resulting in 32,976 training images and 3,000 testing images.
VinDr-SpineXR [73] https://vindr.ai/datasets/spinexr A dataset of annotated medical images for detecting and classifying spinal lesions from radiographs.
NIH ChestX-ray14 [74] https://nihcc.app.box.com/v/ChestXray-NIHCC A dataset comprising 112,120 frontal-view X-ray images from 30,805 unique patients, featuring fourteen disease labels identified through NLP techniques, collected between 1992 and 2015.
CheXpert [58] https://stanfordmlgroup.github.io/competitions/chexpert/ A dataset of chest radiographs from 65,240 patients.
@@
DRD [75] https://www.kaggle.com/competitions/diabetic-retinopathy-detection A dataset consists of high-resolution retina images labeled by subject ID and eye side, with each image rated for diabetic retinopathy severity on a scale of 0 (no DR) to 4 (proliferative DR) by a clinician.
@@
LC25000 (Lung) [61] https://github.com/tampapath/lung_colon_image_set A dataset consists of 15,000 pathology images from lung tissue, including benign tissues and various carcinomas.
@@
MedFMC (Chest) [71] https://github.com/openmedlab/MedFM A dataset for screening thoracic diseases encompasses 19 common thoracic abnormalities.
@@
MedFMC (Endo) [71] https://github.com/openmedlab/MedFM A dataset aimed at enhancing the automatic detection and classification of four different lesion types in colonoscopy images, facilitating early diagnosis of colorectal cancer.
HAM10000 [76] https://www.kaggle.com/datasets/kmader/skin-cancer-mnist-ham10000 This dataset is composed of dermatoscopic images of pigmented lesions.
@@
BUSBRA [77] https://zenodo.org/records/8231412 A dataset of anonymized breast ultrasound images from 1,064 patients, including biopsy-proven tumors, BI-RADS annotations, and ground truth delineations of tumoral and normal regions.
@@
WCE [78] https://www.kaggle.com/datasets/francismon/curated-colon-dataset-for-deep-learning A dataset of colon disease images containing curated samples for training and testing, derived from the Kvasir [79] and ETIS-Larib-Polyp DB [80] datasets.
@@
Fundus JSIEC [81] https://www.kaggle.com/datasets/linchundan/fundusimage1000 A dataset of 1,000 fundus images belonging to 39 classes, sourced from the Joint Shantou International Eye Centre in Shantou city, Guangdong province, China.
@@
HyperKvasir [82] https://datasets.simula.no/hyper-kvasir/ A dataset consisting of 10,662 labeled images in JPEG format, categorized into 23 classes of medical findings, with each class represented by a specific folder, highlighting the imbalance in the number of images per class.
@@
Kvasir [79] https://datasets.simula.no/kvasir/ A dataset of annotated and verified gastrointestinal tract images, featuring various classes of anatomical landmarks and pathological findings, organized in separate folders with resolutions ranging from 720x576 to 1920x1072 pixels, suitable for tasks such as image retrieval and machine learning.
@@
ODIR [83] https://www.kaggle.com/datasets/andrewmvd/ocular-disease-recognition-odir5k A dataset of 5,000 patients featuring age, color fundus photographs from both left and right eyes, along with doctors’ diagnostic keywords, with 6,392 samples used for training.
@@
BUID [84] https://www.kaggle.com/datasets/aryashah2k/breast-ultrasound-images-dataset A dataset of 780 breast ultrasound images from 600 female patients aged 25 to 75, collected in 2018, categorized into three classes: normal, benign, and malignant, with ground truth images included.
@@
PAD-UFES-20 [85] https://www.notion.so/c1e0ff8f36814b94b205ee919e2e3ff8?pvs=21 A dataset of skin lesion images labeled into six categories: Basal Cell Carcinoma, Squamous Cell Carcinoma, Actinic Keratosis, Seborrheic Keratosis, Melanoma, and Nevus.
OCTMNIST [86] https://medmnist.com/ A dataset for multi-class classification of retinal OCT images.
@@
Breast Tumor MRI [87] https://www.kaggle.com/datasets/masoudnickparvar/brain-tumor-mri-dataset A combined dataset (including three datasets: figshare [88], SARTAJ dataset [89]. and Br35H dataset [90]) of 7,023 human brain MRI images classified into four classes: glioma, meningioma, no tumor, and pituitary. The no tumor class images are from the Br35H dataset.
@@
Retinal OCT [91] https://www.kaggle.com/datasets/obulisainaren/retinal-oct-c8 A dataset of high-quality retinal OCT images categorized into 8 classes of retinal diseases, designed for research and model training in classification using machine learning and deep learning.
@@
COVIDxCT [92] https://www.kaggle.com/datasets/hgunraj/covidxct/data A dataset consisting of two variants, ”A” with confirmed COVID-19 cases and ”B” which includes weakly verified cases, designed for training, validation, and testing of CT scans. We choose the “A” set in our experiment.
@@
PMC-9K - This newly curated dataset consists of a cleaned, held-out collection of 9,222 image-text pairs in MedConcept-23M. It is designed to comprehensively evaluate the cross-modal retrieval capabilities of ConceptCLIP across various modalities.
Quilt-1M [16] https://quilt1m.github.io/ This dataset is a large-scale histopathology collection containing 768,826 image-text pairs. For evaluation purposes, we filter out the data sourced from PubMed Central to avoid data leakage and utilize a held-out subset comprising pathology image-text pairs to assess the retrieval performance.
@@
MIMIC-CXR [39] https://physionet.org/content/mimic-cxr/2.1.0/ A dataset contains 371,920 chest X-rays linked to 227,943 imaging studies from 65,079 patients.
IU X-Ray [40] https://www.kaggle.com/datasets/raddar/chest-xrays-indiana-university/data A dataset consists of 7,470 pairs of chest X-ray images and their corresponding diagnostic reports.
@@
VQA-RAD [43] https://osf.io/89kps/ A dataset called VQA-RAD is manually constructed, featuring 3,064 question-answer pairs where clinicians ask natural questions about radiology images and provide reference answers.
SLAKE [42] https://www.med-vqa.com/slake/ A bilingual radiology VQA dataset that includes 642 images and 14,000 questions. We only use the English part.
@@
BRACS-3 [93] https://www.bracs.icar.cnr.it/ A dataset contains 6 different subtypes of lesions including also images representing atypical lesions. Histological images representing normal tissue samples are also included. In this setting, a breast tumor according to a pathology image can be classified into “benign”, “atypical”, or “malignant”.
BRACS-7 [93] https://www.bracs.icar.cnr.it/ A dataset contains 6 different subtypes of lesions including also images representing atypical lesions. Histological images representing normal tissue samples are also included. In this setting, a breast tumor according to a pathology image can be classified into “normal”, “pathological benign”, “usual ductal hyperplasia”, “flat epithelial atypia”, “atypical ductal hyperplasia”, “ductal carcinoma in situ”, or “invasive carcinoma”.
BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset features 985 histopathology whole slide images of Breast Invasive Carcinoma, including 787 Invasive Ductal Carcinoma and 198 Invasive Lobular Carcinoma cases.
NSCLC [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset contains 1,053 histopathology slides of Non-Small Cell Lung Cancer, including 541 lung adenocarcinoma and 512 lung squamous cell carcinoma cases.
Camelyon [64, 95] https://camelyon16.grand-challenge.org/https://camelyon17.grand-challenge.org/ A dataset based on CAMELYON16 [64] and CAMELYON17 [95], which evaluates new and existing algorithms for automated detection and classification of breast cancer metastases in whole-slide images of histological lymph node sections.
Molecular Subtyping
BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset includes 985 histopathology whole slide images of breast invasive carcinoma, specifically 787 Invasive Ductal Carcinoma and 198 Invasive Lobular Carcinoma cases.
@@
BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis. We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
LUAD [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis. We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
LUSC [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis. We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
@@
Dataset Classes Prompts Top-K
SIIM-ACR [59] No Finding, Pneumothorax a chest radiology presents {} 256
RFMiD2 [70] Neovascularization, Macular Edema, Myopia, Retinal Traction, Coloboma, Choroidal folds, Tortuous Vessels, Retinitis Pigmentosa, Retinal pigment epithelium changes, Optic Disc Pallor, Media Haze, Retinitis, Preretinal Hemorrhage, Asteroid Hyalosis, Drusens, Hemorrhagic Pigment Epithelial Detachment, Branch Retinal Vein Occlusion, Optic Disc Edema, Exudation, Haemorrhagic Retinopathy, Tilted Disc, Tessellation, Retinal Tears, Retinal Detachment, Optic Disc Cupping, Macular Hole, Silicone Oil-Filled Eye, Cotton Wool Spots, Vasculitis, Microaneurysm, Macular Scar, Age-Related Macular Degeneration, Optic neuritis, Anterior Ischemic Optic Neuropathy, Laser Scar, Chorioretinitis, Within Normal Limit, Epiretinal Membrane, Central retinal vein occlusion, Central Serous Retinopathy, Optociliary Shunt an image of {} 4
DRD [75] No diabetic retinopathy, Mild diabetic retinopathy, Moderate diabetic retinopathy, Severe diabetic retinopathy, Proliferative diabetic retinopathy a detailed view of a retina indicating {};
@@
32
Covid-CXR2 [60] No Finding, Covid-19 a radiographic representation assessing for {} 256
@@
256
MedFMC (Chest) [71] pleural effusion, nodule, pneumonia, cardiomegaly, hilar enlargement, fracture old, fibrosis, aortic calcification, tortuous aorta, thickened pleura, TB, pneumothorax, emphysema, atelectasis, calcification, pulmonary edema, increased lung markings, elevated diaphragm, consolidation lung situation of {} 2
@@
32
VinDr-Mammo [72] birads negative, breast heterogeneously density, No Finding, breast scattered areas of fibroglandular, birads suspicious malignant, Mass, breast extremely density, birads highly suggestive of malignant, Suspicious Calcification, birads benign, breast almost entirely fatty, Suspicious Lymph Node, Focal Asymmetry, birads probably benign, Asymmetry, Architectural Distortion, Skin Thickening, Global Asymmetry, Nipple Retraction, Skin Retraction an image of {} 32
HAM10000 [76] actinic keratoses and intraepithelial carcinoma, basal cell carcinoma, benign keratosis-like lesions, dermatofibroma, melanoma, melanocytic nevi, vascular lesions a dermatology has that of {} presented 32
VinDr-PCXR [50] Bronchiolitis, Pneumonia, Other disease, Bronchitis, Brocho-pneumonia, Tuberculosis, Mediastinal tumor, nan, Situs inversus, Hyaline membrane disease, CPAM, No finding an image of {} 4

### v2

@@ Discussion
BLEU. The Bilingual Evaluation Understudy (BLEU) score [46] quantifies the similarity between machine-generated text and reference human-authored texts by measuring lexical and syntactic overlap. Originally developed for machine translation, BLEU is widely adopted in medical report generation to evaluate both semantic fidelity (alignment with clinical content) and fluency (linguistic coherence). The score is computed as:
$\text{BLEU}=\underbrace{\text{BP}}_{\text{Brevity Penalty}}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$
where $p_{n}$ represents the modified n-gram precision (the fraction of machine-generated n-grams that match the reference texts), $w_{n}$ are weighting factors (typically $w_{n}=\frac{1}{N}$ for n-grams up to length $N=4$), and BP is the brevity penalty $\min\left(1,e^{1-\frac{\text{reference length}}{\text{generated length}}}\right)$ that prevents artificially high scores for shorter outputs. Higher BLEU scores (range: 0–1) indicate stronger alignment with reference texts. In medical applications, this metric ensures generated reports preserve critical clinical terminology (via $n$-gram precision) while maintaining natural narrative structure (via brevity penalty).
METEOR. The Metric for Evaluation of Translation with Explicit ORdering (METEOR) [47] is another metric for evaluating text generation quality, focusing on precision, recall, and alignment of phrases. It is used for medical report generation to provide a more nuanced assessment of generated text quality.
ROUGE-L. The Recall-Oriented Understudy for Gisting Evaluation (ROUGE-L) [48] measures the longest common subsequence (LCS) between the generated and reference text sequences. It is used for evaluating medical report generation to capture the overlap in content. The formula for ROUGE-L is:
@@
CIDEr. Consensus-based Image Description Evaluation (CIDEr) [49] is a metric designed to assess the similarity of a generated text to multiple reference text sequences, emphasizing consensus. It is used for medical report generation to evaluate the relevance and quality of the generated content.
Clinical Efficacy. Clinical efficacy assesses the practical utility of AI-generated medical reports by measuring their concordance with clinical standards and potential impact on diagnostic decisions. We employ micro-averaged precision, recall, and F1 scores computed using the CheXbert classifier [50], which identifies 14 standardized radiological observations (including pneumothorax and edema) in both generated and reference reports. The micro-averaged approach aggregates performance across all disease classes: Micro Precision measures the fraction of correctly identified findings among all system-predicted abnormalities (true positives/[true positives + false positives]), while Micro Recall calculates the proportion of reference report findings correctly detected by the system (true positives/[true positives + false negatives]). The Micro F1 score, defined as:
[…]
provides a balanced measure of diagnostic performance by combining precision and recall. This comprehensive evaluation framework ensures clinically meaningful assessment of report quality, where high precision minimizes overdiagnosis and high recall reduces missed findings. These metrics reflect system-wide reliability: high precision minimizes overdiagnosis (false positives), while high recall mitigates missed diagnoses (false negatives). Micro-averaging prioritizes per-case accuracy over per-class balancing, mirroring clinical workflows where all observations in a report hold equal urgency.
@@
The pre-training dataset, described in Section Method, is derived from publicly available sources and will be released to the research community upon publication to ensure reproducibility. All evaluation datasets referenced in Extended Data Table LABEL:tab:datasets are publicly accessible benchmarks, with complete citation information provided to enable direct access by other researchers.
@@
The implementation of ConceptCLIP is available at https://github.com/JerrryNie/ConceptCLIP. The weights of ConceptCLIP is released at https://huggingface.co/JerrryNie/ConceptCLIP.
@@
Y.N., S.H., Y.B., and H.C. conceived and designed the work. Y.N., S.H., Y.B., Y.W., Zhixuan C., Zhiyuan C., and H.W. collected the data for pre-training and downstream task evaluation. Y.N. and S.H. contributed to the technical implementation of the pre-training process. Y.B. contributed to the technical implementation of explainable AI. Y.N. and S.H. evaluated medical image diagnosis tasks. Y.N. evaluated cross-modal retrieval tasks. Y.B. evaluated explainable AI tasks. Y.W. evaluated whole-slide image analysis tasks. Zhixuan C. evaluated medical report generation tasks. S.Y. evaluated medical visual question answering tasks. Y.N., Y.B., and S.H. wrote the manuscript with inputs from all authors. Xi.W., L.L., M.W., Xian.W., R.C.K.C., Y.M.L., Y.Z., and P.R. provided suggestions on the framework and experiments. All authors reviewed and approved the final paper. H.C. supervised the research.
@@
Ethics declarations
This study has been reviewed and approved by the Human and Artefacts Research Ethics Committee (HAREC). The protocol number is HREP-2024-0212.
Acknowledgements
This work was supported by the Hong Kong Innovation and Technology Commission (Project No. MHP/002/22, GHP/006/22GD and ITCPD/17-9), HKUST (Project No. FS111), and the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Reference Number: T45-401/22-N),
@@ Appendix A Extended Data
Table 1: Medical image analysis datasets grouped by task. Click the hyperlink in the “link” column to access the corresponding dataset.
Dataset Description Link
@@
SIIM-ACR [51] A binary classification dataset containing chest radiographic images.
Covid-CXR2 [52] A dataset of over 16,000 chest X-ray images from more than 15,100 patients across 51 countries, including 2,300 positive COVID-19 images, designed to differentiate between no pneumonia, non-COVID-19 pneumonia, and COVID-19 pneumonia in the COVIDx V8A dataset, while the COVIDx V8B dataset focuses on detecting COVID-19 positive and negative cases.
NLM-TB [26] A radiology image dataset with 800 total images with the label of normal or tuberculosis.
LC25000 (Colon) [53] A dataset comprises 10,000 pathology images from colon tissues, featuring benign tissues and adenocarcinomas.
UBIBC [54] A dataset consists of ultrasound images related to breast cancer, including benign and malignant.
PCam200 [55] A dataset of pathological H&E images created from the Camelyon2016 challenge dataset [56].
RSNA [43] A dataset of about 30,000 frontal view chest X-ray images, each labeled for binary classification to indicate the presence of pneumonia.
Brain Tumor CT [57] A dataset of CT scans for brain tumor detection and analysis. It features high-resolution images from multiple patients, labeled with tumor types (e.g., glioma, meningioma) and their locations in the brain, designed to aid in developing AI models for automatic detection, classification, and segmentation of brain tumors.
Brain Tumor MRI [57] This dataset is similar to the Brain Tumor CT dataset [57] but with the MRI modality.
DDSM [58] A dataset containing 55,890 pre-processed images from the DDSM [59] and CBIS-DDSM [60] datasets, formatted as 299$\times$299 pixel TFRecords for TensorFlow, with 14% positive and 86% negative examples, split into test and validation sets.
Breast Cancer [61] A dataset of 3,383 annotated mammogram images focused on breast tumors, exported from Roboflow, designed for building and testing deep learning models for tumor detection.
HMC-QU [62] A dataset created by Hamad Medical Corporation, Tampere University, and Qatar University includes 2D echocardiography recordings from 2018 and 2019 for heart attack detection and left ventricle wall segmentation, with recordings at 25 fps and resolutions from 422$\times$636 to 768$\times$1024 pixels.
@@
RFMiD2 [63] A multi-label dataset of fundus images annotated by three eye specialists.
MedFMC (Colon) [64] A dataset focused on facilitating the detection of early-stage cancer cells in tissue slides, enabling pathologists to classify and quantify cancerous regions.
VinDr-Mammo [65] A large-scale benchmark dataset for computer-aided detection and diagnosis in full-field digital mammography.
VinDr-PCXR [45] A dataset focuses on pediatric chest X-rays for the interpretation of common thoracic diseases.
VinDr-CXR [44] A dataset of chest X-Ray images containing 18,000 high-quality postero-anterior scans with detailed localization of 22 critical findings and classification of 6 thoracic diseases, annotated by experienced radiologists from major hospitals in Vietnam.
VinDr-SpineXR [66] A dataset of annotated medical images for detecting and classifying spinal lesions from radiographs.
NIH ChestX-ray14 [67] A dataset comprising 112,120 frontal-view X-ray images from 30,805 unique patients, featuring fourteen disease labels identified through natural language processing (NLP) techniques, collected between 1992 and 2015.
CheXpert [50] A dataset of chest radiographs from 65,240 patients.
@@
DRD [68] A dataset consists of high-resolution retina images labeled by subject ID and eye side, with each image rated for diabetic retinopathy severity on a scale of 0 (no DR) to 4 (proliferative DR) by ophthalmologists.
@@
LC25000 (Lung) [53] A dataset consists of 15,000 pathology images from lung tissue, including benign tissues and various carcinomas.
@@
MedFMC (Chest) [64] A dataset for screening thoracic diseases encompasses 19 common thoracic abnormalities.
@@
MedFMC (Endo) [64] A dataset aimed at enhancing the automatic detection and classification of four different lesion types in colonoscopy images, facilitating early diagnosis of colorectal cancer.
HAM10000 [69] A dataset consists of dermatoscopic images of pigmented lesions.
@@
BUSBRA [70] A dataset of anonymized breast ultrasound images from 1,064 patients, including biopsy-proven tumors, BI-RADS annotations, and ground truth delineations of tumoral and normal regions.
@@
WCE [71] A dataset of colon disease images containing curated samples for training and testing, derived from the Kvasir [72] and ETIS-Larib-Polyp DB [73] datasets.
@@
Fundus JSIEC [74] A dataset of 1,000 fundus images belonging to 39 classes, sourced from the Joint Shantou International Eye Centre in Shantou city, Guangdong province, China.
@@
HyperKvasir [75] A dataset consisting of 10,662 labeled images in the JPEG format, categorized into 23 classes of medical findings, highlighting the imbalance in the number of images per class.
@@
Kvasir [72] A dataset of annotated and verified gastrointestinal tract images, featuring various classes of anatomical landmarks and pathological findings, with resolutions ranging from 720$\times$576 to 1920$\times$1072 pixels.
@@
ODIR [76] A dataset of 5,000 patients featuring color fundus photographs from both left and right eyes, along with ophthalmologists’ diagnostic keywords, with 6,392 samples used for training.
@@
BUID [77] A dataset of 780 breast ultrasound images from 600 female patients aged 25 to 75, collected in 2018, categorized into three classes: normal, benign, and malignant.
@@
PAD-UFES-20 [78] A dataset of skin lesion images labeled into six categories: basal cell carcinoma, squamous cell carcinoma, actinic keratosis, seborrheic keratosis, melanoma, and nevus.
OCTMNIST [79] A dataset for multi-class classification of retinal OCT images.
@@
Breast Tumor MRI [80] A combined dataset (including three datasets: figshare [81], SARTAJ dataset [82]. and Br35H dataset [83]) of 7,023 human brain MRI images classified into four classes: glioma, meningioma, no tumor, and pituitary. The “no tumor” class images are from the Br35H dataset.
@@
Retinal OCT [84] A dataset of high-quality retinal OCT images categorized into 8 classes of retinal diseases.
@@
COVIDxCT [85] A dataset of CT scans consisting of two variants, “A” with confirmed COVID-19 cases and “B” which includes weakly verified cases. We choose the “A” set in our experiment.
@@
PMC-9K newly curated dataset consisting of a cleaned, held-out collection of 9,222 image-text pairs in MedConcept-23M. It is designed to comprehensively evaluate the cross-modal retrieval capabilities of ConceptCLIP across various modalities. -
Quilt-1M [16] A large-scale histopathology collection containing 768,826 image-text pairs. For evaluation purposes, we filter out the data sourced from PubMed Central and utilize a held-out subset comprising pathology image-text pairs to assess the retrieval performance.
@@
MIMIC-CXR [17] A dataset contains 371,920 chest X-rays linked to 227,943 imaging studies from 65,079 patients.
IU X-Ray [18] A dataset consists of 7,470 pairs of chest X-ray images and their corresponding diagnostic reports.
@@
VQA-RAD [38] A dataset manually constructed, featuring 3,064 question-answer pairs where clinicians ask questions about radiology images and provide reference answers.
SLAKE [37] A bilingual radiology VQA dataset that includes 642 images and 14,000 questions. We only use the English part.
@@
BRACS-3 [86] A dataset contains 6 different subtypes of lesions, including images representing atypical lesions. Histological images representing normal tissue samples are also included. In this setting, a breast tumor according to a pathology image can be classified into “benign”, “atypical”, or “malignant”.
BRACS-7 [86] A dataset that contains the same images as BRACS-3 but with a different label set. In this setting, a breast tumor according to a pathology image can be classified into “normal”, “pathological benign”, “usual ductal hyperplasia”, “flat epithelial atypia”, “atypical ductal hyperplasia”, “ductal carcinoma in situ”, or “invasive carcinoma”.
NSCLC [87] A dataset contains 1,053 histopathology slides of non-small cell lung cancer, including 541 lung adenocarcinoma and 512 lung squamous cell carcinoma cases.
Camelyon [56, 88] A dataset based on CAMELYON16 [56] and CAMELYON17 [88], which evaluates new and existing algorithms for automated detection and classification of breast cancer metastases in whole-slide images of histological lymph node sections.
Mutation Prediction
BRCA-TP53 [87] This dataset comprises molecular and clinical data from the Breast Invasive Carcinoma (BRCA) cohort of The Cancer Genome Atlas (TCGA), with specialized annotation for TP53-mutant tumors, the most frequently altered tumor suppressor in this malignancy.
BRCA-TNN [87] This dataset comprises molecular and clinical profiles of triple-negative breast tumors (TNBC) from the TCGA BRCA cohort, defined by immunohistochemical absence of estrogen receptor (ER), progesterone receptor (PR), and HER2 amplification.
LUAD-TP53 [87] This dataset provides comprehensive molecular profiling of TP53-mutant lung adenocarcinoma cases from the TCGA cohort.
@@
BRCA [87] A TCGA dataset of breast carcinoma for survival prediction.
LUSC [87] A dataset of the Lung Squamous Cell Carcinoma LUSC cohort from the TCGA for survival prediction.
@@
Dataset Classes Prompts
SIIM-ACR [51] No Finding, Pneumothorax a chest radiology presents {}
RFMiD2 [63] Neovascularization, Macular Edema, Myopia, Retinal Traction, Coloboma, Choroidal Folds, Tortuous Vessels, Retinitis Pigmentosa, Retinal Pigment Epithelium Changes, Optic Disc Pallor, Media Haze, Retinitis, Preretinal Hemorrhage, Asteroid Hyalosis, Drusens, Hemorrhagic Pigment Epithelial Detachment, Branch Retinal Vein Occlusion, Optic Disc Edema, Exudation, Haemorrhagic Retinopathy, Tilted Disc, Tessellation, Retinal Tears, Retinal Detachment, Optic Disc Cupping, Macular Hole, Silicone Oil-Filled Eye, Cotton Wool Spots, Vasculitis, Microaneurysm, Macular Scar, Age-Related Macular Degeneration, Optic Neuritis, Anterior Ischemic Optic Neuropathy, Laser Scar, Chorioretinitis, Within Normal Limit, Epiretinal Membrane, Central Retinal Vein Occlusion, Central Serous Retinopathy, Optociliary Shunt an image of {}
DRD [68] No Diabetic Eetinopathy, Mild Diabetic Retinopathy, Moderate Diabetic Retinopathy, Severe Diabetic Retinopathy, Proliferative Diabetic Retinopathy a detailed view of a retina indicating {};
@@
COVID-CXR2 [52] No Finding, COVID-19 a radiographic representation assessing for {}
@@
MedFMC (Chest) [64] Pleural Effusion, Nodule, Pneumonia, Cardiomegaly, Hilar Enlargement, Fracture Old, Fibrosis, Aortic Calcification, Tortuous Aorta, Thickened Pleura, TB, Pneumothorax, Emphysema, Atelectasis, Calcification, Pulmonary Edema, Increased Lung Markings, Elevated Diaphragm, Consolidation lung situation of {}
@@
VinDr-Mammo [65] Birads Negative, Breast Heterogeneously Density, No Finding, Breast Scattered Areas of Fibroglandular, Birads Suspicious Malignant, Mass, Breast Extremely Density, Birads Highly Suggestive of Malignant, Suspicious Calcification, Birads Benign, Breast Almost Entirely Fatty, Suspicious Lymph Node, Focal Asymmetry, Birads Probably Benign, Asymmetry, Architectural Distortion, Skin Thickening, Global Asymmetry, Nipple Retraction, Skin Retraction an image of {}
HAM10000 [69] Actinic Keratoses and Intraepithelial Carcinoma, Basal Cell Carcinoma, Benign Keratosis-Like Lesions, Dermatofibroma, Melanoma, Melanocytic Nevi, Vascular Lesions a dermatology has that of {} presented
VinDr-PCXR [45] Bronchiolitis, Pneumonia, Other Disease, Bronchitis, Brocho-Pneumonia, Tuberculosis, Mediastinal Tumor, Nan, Situs Inversus, Hyaline Membrane Disease, CPAM, No Finding an image of {}